If a heating coil of the R=2.5Ω is switched on with a I = 5A for 8 minutes to heat up a 2 kg liquid, its temperature rises from T= 20oC to 55oC. If we assume that there is no phase change, no heat loss and the heat absorbed by the heating coil is 280J, what is the liquid specific heat capacity C (J/kg/K))?

Select one:

a. 428.6

b. 3.14

c. 270.2

d. 424.6
Answer #1

Power generated by coil = i2 * Resistance * time(s)

power = 5^2 * 2.5 * 8 * 60

power = 30000 Joule

now coil absorb 280 J of energy so energy left for liquid = 30000 - 280

E = 29720 Joule

now this enegy will be equal to melting energy required for liquid

so E = m * C * delta_T

29720 = 2 * C * 35

C = 424.57   

C = 424.6 (j/kg-k) <<<<<<<<<OPTION D